Jane's utility function is represented as: U = F0.5 C0.5, F is quantity of food and C is quantity of clothing. If her budget constraint is represented as: 125 = 2F + 4C, her optimal bundle of consumption contains how many units of clothing?

the answer is 15.6250 please show steps
Answer #1

For optimal consumption/utility maximization:

MUf/Pf = MUc/Pc

MUf = dU/dF = 0.5C0.5/F0.5

MUc = dU/dC= 0.5F0.5/C0.5

Pf = 2 , Pc = 4

Therefore, MUf/Pf = MUc/Pc would yield:

C/F = 1/2

Replacing F = 2C in the budget constraint:

125 = 2C x 2 + 4C

C = 125/8 = 15.625